OpenStudy (anonymous):

2x - 3y = 7 and x - 4y = 2 what is the first step in the given step in solving the given system of equations by the addition method

OpenStudy (anonymous):

distributing 2?

OpenStudy (anonymous):

But what is the first step for the first one and second one

OpenStudy (anonymous):

if u are using the addition method u keep the first equation the same and them distribute -2 to the second x - 4y = 2 changes to -2x+8y=-4 them you add the two functions

OpenStudy (anonymous):

2x - 3y = 7 +++++ -2x +8y = -4====== 5y+3

OpenStudy (anonymous):

2x - 3y = 7 +++++ -2x +8y = -4====== 5y=3

OpenStudy (anonymous):

ok thxs